You live near Mac Donnel's and want to estimate the proportion ,p, of your neighbors who go there on a typical day.

a. How many people in your neighbors should include in your random sampling to be 97% sure that you are within 7% of p?

b. If 21 out of 25 of Ashley's neighbors visit Mac Donnel's every day, how many people should include in your random sampling to be 86% sure that you are within 3% of p?

********CAN YOU PLEASE SHOW ME HOW TO DO THIS ON A TI-83 PLUS CALCULATOR ********
Answer #1

Ans:

Use invNorm function of TI-84,to get critical z value:

invNorm(alpha/2,0,1)

a)alpha=1-0.97=0.03

alpha/2=0.015

z value=2.17 for 97% CI

Margin of error,E=0.07

Sample size required,n=2.17^2*0.5*(1-0.5)/0.07^2=240

b)prior estimate of p=21/25=0.84

z=1.476

Margin of error,E=0.03

Sample size required,n=1.476^2*0.84*(1-0.84)/0.03^2=325